OpenStudy (anonymous):

8e^3x+9=13 round to the nearest tenth? its logarithims -_-....

OpenStudy (anonymous):

\[8e^{3x}=4\]

OpenStudy (anonymous):

\[e^{3x}=\frac{1}{2}\]

OpenStudy (anonymous):

oops i meant thousandths.... and it has to be in decimal form.

OpenStudy (anonymous):

now what can you do to get the e to cancel

OpenStudy (anonymous):

divide it by itself??

OpenStudy (anonymous):

\[\ln(e^x)=x\]

OpenStudy (anonymous):

that is basically all you need to remember whe you want to get rid of an e raised to variable

OpenStudy (anonymous):

so if you multiply both sides by ln

OpenStudy (anonymous):

\[\ln(e^{3x})=\ln(\frac{1}{2})\] \[3x=\ln(\frac{1}{2})\]

OpenStudy (anonymous):

\[\ln(\frac{a}{b})=\ln(a)-\ln(b) \] this is the next rule you'll need

OpenStudy (anonymous):

\[3x=\ln(1)-\ln(2)\] \[x=\frac{\ln(1)}{3}-\frac{\ln(2)}{3}\]

OpenStudy (anonymous):

\[x=\frac{\ln(1)-\ln(2)}{3}\]

OpenStudy (anonymous):

-0.231?

OpenStudy (anonymous):

if thats what you got then yes